OpenStudy (anonymous):

convert 2sqrt3+2i into its polar form

OpenStudy (anonymous):

@mathstudent55 can you help???

zepdrix (zepdrix):

Hey there Jules c: \[\Large\rm a+b \mathcal i\] Converting to polar form:\[\Large\rm r(\cos \theta+\mathcal i \sin \theta)\]Where r is given by,\[\Large\rm r=\sqrt{a^2+b^2}\]and the angle theta is found by using,\[\Large\rm \tan \theta=\frac{b}{a}\]

zepdrix (zepdrix):

So for our problem here,\[\Large\rm 2\sqrt3+2\mathcal i\]Our r value,\[\Large\rm r=\sqrt{(2\sqrt3)^2+(2)^2}\]Simplify to get your radius. And our theta,\[\Large\rm \tan \theta=\frac{2}{2\sqrt3}=\frac{1}{\sqrt3}\]This corresponds to one of your special angles. Since both the `real` and `imaginary` components of our point were `positive`, we're in the first quadrant.